# DP-BREM: Differentially-Private and Byzantine-Robust Federated Learning with Client Momentum

### v2

@@ 1 Introduction
[…]
DP-BREM+ (our final solution) $\dagger$ untrusted clients (jointly)
[…]
    We show the privacy-utility tradeoff by fixing the same privacy cost (in terms of DP) and then comparing the standard deviation of the noise on the aggregation, where a smaller standard deviation means the DP noise has less negative impact on the utility. Note that different approaches have different aggregation strategies, where $g_{i}$ denotes local gradient, and $m_{i}$ denotes local momentum.
  • $\diamond$
    DDP-RP assumes an honest-but-curious server, i.e., following protocol instructions honestly, but may try to learn additional information. It guarantees distributed DP (DDP) with secure aggregation techniques, where clients add partial noise with a smaller standard deviation, depending on the number of honest clients or its lower bound, denoted by $\tau$. Thus, it provides better privacy-utility tradeoff than local DP (LDP).
  • $\dagger$
    DP-BREM+ achieves the same DP and robustness guarantees as DP-BREM, but has a different trust assumption on the server. It achieves the same noise as central DP, i.e., the DP noise is added to the aggregation, but does not require a trusted server because the noise is securely generated by clients (via the proposed noise generation protocol) and securely added to the aggregation (via a secure aggregation protocol).
1) We propose a novel differentially private and Byzantine-robust FL protocol called DP-BREM, which adds DP noise to the aggregated client momentum with carefully computed sensitivity. Our privacy analysis (shown in Theorem 1) accounts for the privacy cost from momentum, which is different from the conventional DP-SGD that accounts for the privacy cost from the gradient. We also provide the convergence analysis of DP-BREM (shown in Theorem 3), which indicates that there is only a small sacrifice in the convergence rate to satisfy DP (compared to the large sacrifice in convergence of the baseline solution shown in Section 3.2).
2) Considering that DP-BREM is developed under the assumption of a trusted server, we propose the final solution called DP-BREM+ (in Section 5), which achieves the same privacy and robustness properties as DP-BREM, even under a malicious server (for privacy only), using secure multiparty computation techniques. DP-BREM+ is built based on the framework of secure aggregation with verifiable inputs (SAVI) [35], but extends it to guarantee the integrity of DP noise via a novel secure distributed noise generation protocol. Our extended SAVI protocol is general enough to be applied to other DP and robust FL protocols that are average-based.
3) We conduct extensive experiments using MNIST and CIFAR-10 datasets (in Section 6) to demonstrate the effectiveness of our protocols. The results show that it can achieve better utility under the same record-level DP guarantees, as well as strong robustness against Byzantine clients under state-of-the-art attacks, compared to the baseline methods.
@@ 2.1 Differential Privacy (DP)
DP-SGD Algorithm. The most well-known differentially-private algorithm in machine learning is DP-SGD [1], which introduces two modifications to the vanilla stochastic gradient descent (SGD). First, a clipping step is applied to the gradient so that the gradient is in effect bounded for a finite sensitivity. The second modification is Gaussian noise augmentation on the summation of clipped gradients, which is equivalent to applying the Gaussian mechanism to the updated iterates. The privacy accountant of DP-SGD is shown in Appendix F.
@@ 2.2 Federated Learning (FL) with DP
Federated Learning (FL) [21, 29] is a collaborative learning setting to train machine learning models. We consider the horizontal cross-silo FL setting, which involves multiple clients, each holding their own private dataset of the same set of features, and a central server that implements the aggregation. Unlike the traditional centralized approach, data is not stored at a central server; instead, clients train models locally and exchange updated parameters with the server, which aggregates the received local model parameters and sends them to the clients. Based on the participating clients and scale, federated learning can be classified into two types: cross-device FL where clients are typically mobile devices and the client number can reach up to a scale of millions; cross-silo FL (our focus) where clients are organizations or companies and the client number is usually small (e.g., within a hundred).
FL with DP. In FL, the neighboring datasets $D$ and $D^{\prime}$ in Definition 1 can be defined at two distinct levels: record-level and client-level. In cross-device FL, each device usually stores one individual’s data, then the whole devices’ data should be protected. It corresponds to client-level DP, where $D^{\prime}$ is obtained by adding or removing one client/device’s whole training dataset from $D$. In cross-silo FL, each record corresponds to one individual’s data, then record-level DP should be provided, where $D^{\prime}$ is obtained by adding or removing a single training record/example from $D$. Since we consider cross-silo FL, achieving record-level DP is our privacy goal.
@@ 3.2 Challenges and Baseline
Challenges: replacing average-based aggregator leads to large sensitivity of DP. Though there are many works on achieving either DP or Byzantine robustness, it’s nontrivial to achieve both with high utility. The main reason is that most Byzantine-robust methods replace the averaging aggregator with median-based strategies or some complex robust aggregators, which leads to a large sensitivity of DP compared to the averaging operation, as illustrated in Example 1.
@@
DP-LFH: baseline via direct combination of LFH and DP-SGD. As shown in Section 2.3, the Byzantine-robust scheme LFH [22] utilizes an average-based aggregator, which can be regarded as a non-private robust solution to address the disadvantage of median-based aggregator. A straightforward method to add DP protection on top of LFH is to combine it with the DP-SGD algorithm. However, LFH requires each client to compute the local momentum $\bm{m}_{t,i}$ for server aggregation, while DP-SGD aggregates gradients and accounts for the privacy cost via the composition of iterative gradient update. Without a trusted server, a straightforward solution to combine DP with LFH is to use DP-SGD at each client to privatize the local gradient, and then compute the momentum from the privatized gradient (thus there is no additional privacy cost due to post-processing). Formally, client $\mathsf{C}_{i}$ computes
$\bm{g}_{t,i}=\frac{1}{|\mathcal{D}_{i,t}|}\left[\sum\nolimits_{\bm{x}\in\mathcal{D}_{i,t}}\mathsf{Clip}_{R}(\nabla_{\bm{\theta}}\ell(\bm{x},\bm{\theta}_{t-1}))+\mathcal{N}(0,R^{2}\sigma^{2}\mathbf{I}_{d})\right],$ (7)
where $\mathbf{I}_{d}$ is an identity matrix with size $d\times d$ ($d$ is the model size, i.e., $\bm{\theta}_{t}\in\mathbb{R}^{d}$), the record-level clipping parameter $\mathsf{Clip}_{R}(\cdot)$ restricts the sensitivity when adding/removing one record from the local dataset, and Gaussian noise $\mathcal{N}(0,R^{2}\sigma^{2}\mathbf{I}_{d})$ introduces DP property on $\bm{g}_{t,i}$. Since DP is immune to post-processing, the remaining steps can be implemented in the same way as the original LFH, without incurring additional privacy costs. This baseline solution DP-LFH achieves record-level DP against an untrusted server. However, it has several limitations, which lead to both poor privacy-utility tradeoff and robustness.
Limitation 1: large aggregated noise. Since each client locally adds DP noise, the overall noise after aggregation will be larger than the case of the central setting under the same privacy budget (i.e., the value of $\epsilon$), because only the server adds noise in the central setting. Therefore, DP-LFH has a poor privacy-utility tradeoff.
@@
where $\bm{g}_{1,i},\bm{g}_{2,i},\cdots,\bm{g}_{t,i}$ are already privatized via local noise. Assume the server samples a subset of clients for aggregation in each iteration. Assume that client $\mathsf{C}_{i}$’s momentum $\bm{m}_{t,i}$ is not selected in the $t$-th iteration, thus the aggregate is independent of $\bm{g}_{t,i}$. However, in a later iteration (i.e., $\tau>t$), if client $\mathsf{C}_{i}$’s momentum $\bm{m}_{\tau,i}$ is involved in the aggregation, it will depend on $\bm{g}_{t,i}$ according to (9). Therefore, we need to account for the privacy cost of $\bm{g}_{t,i}$ in all iterations. There is no privacy amplification benefit from sampling clients, leading to high privacy costs or low utility.
@@ 4 DP-BREM
To address the limitations of DP-LFH, we start from the assumption of a trusted server that can obtain clients’ local models/updates and generate DP noise, and propose an initial solution called DP-BREM (in Section 4.1). It is a differentially-private version of LFH with carefully designed enhancements, achieving a similar level of robustness as the non-private LFH. Since DP-BREM adds DP noise to the momentum (as versus adding noise to the gradient in DP-SGD), our privacy accountant shown in Section 4.2 is different from the traditional privacy accountant of DP-SGD. We also provide the convergence analysis in Section 4.3, where the provable convergence of LFH is maintained with only a small difference. Based on DP-BREM, we then relax the server’s trust assumption in our final solution DP-BREM+ (in Section 5), by adopting secure multiparty computation techniques including secure aggregation with input validation and joint noise generation, which achieves the same DP guarantee with the same amount of noise as in DP-BREM, without trusting the server.
@@ 4.1 Algorithm Design
The mathematical notations involved in our algorithm design and theoretical analysis are summarized in Table 5 (see Appendix A). The illustration of our design is shown in Figure 1, and the algorithm is shown in Algorithm 1, where all clients need to implement local updates (in Line-3), but only a subset of their momentum vectors are aggregated by the server (in Line-4). The details of client updates and server aggregation are described below.
@@
where $\mathsf{Clip}_{R}(\cdot)$ is the clipping function defined in (4), but is used here to bound the sensitivity for DP (refer to DP-SGD discussed in Section 2.1). Note that the batch size $|\mathcal{D}_{i,t}|$ is random and $\mathbb{E}[|\mathcal{D}_{i,t}|]=p_{i}|\mathcal{D}_{i}|$. Then, the local momentum can be computed by
@@
0:  Initialization $\bm{\theta}_{0}\in\mathbb{R}^{d}$, clipping bounds $R$ and $C$ , learning rate $\eta_{t}$ of the global model.
@@ 4.2 Privacy Analysis
(Sketch) According to (10), the sensitivity of $\bar{\bm{g}}_{t,i}$ is $\frac{R}{p_{i}|\mathcal{D}_{i}|}$ because each clipped term $\mathsf{Clip}_{R}(\cdot)$ has bounded L2-norm, i.e., $\|\mathsf{Clip}_{R}(\cdot)\|\leqslant R$. Then, due to the recursive representation of local momentum in (11), the sensitivity of $\bm{m}_{t,i}$ is $\frac{R}{p_{i}|\mathcal{D}_{i}|}$. Finally, the client-level clipping $\mathsf{Clip}_{C}(\cdot)$ introduces another upper bound for the sensitivity. Refer to Appendix A for the full-version proof. ∎
@@
The final privacy analysis of DP-BREM is shown in Theorem 1. It presents how to compute the privacy budget $\epsilon$ and privacy parameter $\delta$ when the parameters (such as $T$, $\sigma$, $q$, etc.) of the algorithm are given. We use an advanced privacy accountant tool called Gaussian DP (GDP) [11] (refer to Appendix F), then convert it to $(\epsilon,\delta)$-DP. Note that in our privacy analysis, clients can use different record-level sampling rates $p_{i}$, thus different sensitivity $S_{i}$ shown in (14). Therefore, the final privacy budget (denoted by $\epsilon_{i}$) of DP-BREM may be different for different clients, which provides personalized privacy if these parameters are different for each client.
@@
This result is obtained by the composition of multiple iterations and the privacy amplification from sampling. See Appendix B for the detailed proof. ∎
@@ 4.3 Convergence Analysis
After some parameter tuning (the detailed tuning is shown under (C) in Appendix C) of the clipping bounds:
@@
where $|\mathcal{B}|$ is the number of Byzantine clients, $d$ is the dimension of model parameter $\bm{\theta}_{t}$, $\sigma$ is the noise multiplier (for DP) shown in (12), $q$ is the client-level sampling rate shown in Line-2 of Algorithm 1, and $\rho$ is defined in (17). The formal version of (19) is shown in (C) of Appendix C.
@@
(Sketch) Directly bounding $\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{m}_{t}^{*}\|^{2}$ is not easy, thus we utilize the upper bounds of $\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{\mu}\|^{2}$ and $\mathbb{E}\|\bm{\mu}-\bm{m}_{t}^{*}\|^{2}$ to get the final result, where $\bm{\mu}\coloneqq\mathbb{E}[\bm{m}_{t,i}]$ is the expected local momentum (we assume clients’ local momentum are i.i.d.). When upper bounding $\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{\mu}\|^{2}$, we can decompose it to three types of errors: error of honest clients (due to randomness and bias introduced by clipping), error of Byzantine clients (due to Byzantine perturbation), and error introduced by the added DP noise. Furthermore, we can have the optimized parameter tuning of $C$ and $R$ to minimize the summation of the above three types of errors. Refer to Appendix C for the full-version proof. ∎
@@
Table 2: Comparison of Convergence Rate
Where to add noise Convergence Rate
LFH [22] None $O(\rho\sqrt{1+|\mathcal{B}|})$
@@ 5.1 Challenges
Secure Aggregation with Verified Inputs (SAVI). The key crypto technique we leverage to achieve the above objectives is SAVI [35], which is a type of protocols that securely aggregate only well-formed inputs. The security goals include both privacy and integrity. Specifically, privacy means that no party should be able to learn anything about the raw input of an honest client, other than what can be learned from the final aggregation result. Integrity means that the output of the protocol returns the correct aggregate of well-formed input, where 1) an input $u$ passes the input integrity check with a public validation predicate $\mathsf{Valid}(\cdot)$ if and only if $\mathsf{Valid}(u)=1$, and 2) the aggregation is correctly computed. An instantiation of the SAVI protocol is EIFFeL [35] (described in Appendix G).
@@
A straightforward solution is to follow [36] that assumes the existence of another semi-honest server (but does not collude with the original server) that will generate DP noise and execute the privacy engine. However, the assumption of another non-colluding server may not be practical and we assume only a single server.
Another alternative solution is to leverage Distributed DP (DDP) [39], where Gaussian noise is generated by clients in a distributed way: each client generates a Gaussian noise locally, and the aggregation of Gaussian noise also follows a Gaussian distribution with an enlarged standard deviation. Since only the aggregated result is released (with the help of crypto techniques), each client can add a smaller noise with the guarantee that the aggregated noise satisfies the required DP. However, this solution has two limitations in our scenario. First, distributed noise generation needs to add more noise to achieve the same privacy compared with server-side noise generation due to the collusion of malicious clients. Second, malicious clients can generate arbitrary values as the local Gaussian noise, which has a large impact on the robustness.
@@
Overview of DP-BREM+. To achieve secure aggregation with verified inputs and secure Gaussian noise generation under the threat model of a malicious server and malicious minority of clients, our DP-BREM+ 1) leverages an existing SAVI protocol called EIFFeL [35] to achieve secure input validation; and 2) introduces a new protocol to achieve secure noise generation that is compatible with EIFFeL. The idea of jointly generating Gaussian noise in DP-BREM+ is inspired by [34], but our design is based on Shamir’s secret sharing [37] with robust reconstruction by following the design in EIFFeL, thus guarantees security under malicious minority. We present the preliminaries of Shamir’s secret sharing and EIFFeL protocol in Appendix G.

### v3

@@ 1 Introduction
1) We propose DP-BREM, a novel differentially private and Byzantine-robust FL protocol that adds DP noise to aggregated client momentum with computed sensitivity. Our privacy analysis (Theorem 1) accounts for momentum, differing from conventional DP-SGD which accounts for gradient. Our convergence analysis (Theorem 3) shows minimal convergence rate sacrifice for DP compared to the baseline.
2) Considering that DP-BREM is developed under the assumption of a trusted server, we propose the final solution called DP-BREM+ (Section 5), which achieves the same privacy and robustness properties as DP-BREM, even under a malicious server (for privacy only), using secure multiparty computation techniques. DP-BREM+ is built based on the framework of secure aggregation with verifiable inputs (SAVI) [36], but extends it to guarantee the integrity of DP noise via a novel secure distributed noise generation protocol. Our extended SAVI protocol is general enough to be applied to other DP and robust FL protocols that are average-based.
3) We demonstrate our protocols’ effectiveness through extensive experiments on MNIST, CIFAR-10, and FEMNIST datasets (Section 6), showing improved utility with the same record-level DP guarantees and strong robustness against Byzantine clients under state-of-the-art attacks, compared to baseline methods.
@@ 2.1 Differential Privacy (DP)
DP-SGD Algorithm. The most well-known differentially-private algorithm in machine learning is DP-SGD [1], which introduces two modifications to the vanilla stochastic gradient descent (SGD). First, a clipping step is applied to the gradient so that the gradient is in effect bounded for a finite sensitivity. The second modification is Gaussian noise augmentation on the summation of clipped gradients, which is equivalent to applying the Gaussian mechanism to the updated iterates. The privacy accountant of DP-SGD is shown in Appendix I.
@@ 2.2 Federated Learning (FL) with DP
Federated Learning (FL) [22, 30] is a collaborative learning setting to train machine learning models. We consider the horizontal cross-silo FL setting, which involves multiple clients, each holding their own private dataset of the same set of features, and a central server that implements the aggregation. Unlike the traditional centralized approach, data is not stored at a central server; instead, clients train models locally and exchange updated parameters with the server, which aggregates the received local model parameters and sends them to the clients. Based on the participating clients and scale, federated learning can be classified into two types: cross-device FL where clients are typically mobile devices and the client number can reach up to a scale of millions; cross-silo FL (our focus) where clients are organizations or companies and the client number is relatively small (e.g., within hundreds).
FL with DP. In FL, the neighboring datasets $D$ and $D^{\prime}$ in Definition 1 can be defined at two distinct levels: record-level and client-level. In cross-device FL, each device usually stores one individual’s data, and then the whole device’s data should be protected. It corresponds to client-level DP, where $D^{\prime}$ is obtained by adding or removing one client/device’s whole training dataset from $D$. In cross-silo FL, each record corresponds to one individual’s data, then record-level DP should be provided, where $D^{\prime}$ is obtained by adding or removing a single training record/example from $D$. Since we consider cross-silo FL, achieving record-level DP is our privacy goal.
@@ 3.1 Problem Statement
[…]
DP-BREM+ (our final solution) $\dagger$ untrusted clients (jointly)
[…]
    We demonstrate the privacy-utility tradeoff by comparing the standard deviation of DP noise on the aggregation, with smaller values indicating less negative impact on utility. Note that different approaches use different aggregation strategies, where $g_{i}$ is the local gradient and $m_{i}$ is the local momentum.
  • $\diamond$
    DDP-RP assumes an honest-but-curious server and ensures distributed DP (DDP) with secure aggregation. Clients add partial noise with a smaller standard deviation based on the number of honest clients, $\tau$, resulting in a better privacy-utility tradeoff than local DP (LDP).
  • $\dagger$
    DP-BREM+ matches DP-BREM’s DP and robustness guarantees with a different server trust assumption. It achieves central DP without a trusted server, as clients securely generate and add noise using the proposed noise generation and secure aggregation protocols.
@@ 3.2 Challenges and Baseline
Challenges. Replacing the average-based aggregator with median-based or complex robust aggregators increases DP sensitivity. Achieving both DP and Byzantine robustness with high utility is challenging because these methods result in significantly larger DP sensitivity than averaging, as illustrated in Example 1.
@@
DP-LFH: baseline via direct combination of LFH and DP-SGD. As shown in Section 2.3, the Byzantine-robust scheme LFH [23] utilizes an average-based aggregator, which can be regarded as a non-private robust solution to address the disadvantage of the median-based aggregator. A straightforward method to add DP protection on top of LFH is to combine it with the DP-SGD algorithm. However, LFH requires each client to compute the local momentum $\bm{m}_{t,i}$ for server aggregation, while DP-SGD aggregates gradients and accounts for the privacy cost via the composition of iterative gradient update. In LFH, since the gradient is computed only on the client-side, a straightforward solution to integrate DP is to use DP-SGD at each client to privatize the local gradient, and then compute the momentum from the privatized gradient (thus there is no additional privacy cost due to post-processing). Formally, client $\mathsf{C}_{i}$ computes
$\bm{g}_{t,i}=\frac{1}{|\mathcal{D}_{i,t}|}\sum_{\bm{x}\in\mathcal{D}_{i,t}}\mathsf{Clip}_{R}(\nabla_{\bm{\theta}}\ell(\bm{x},\bm{\theta}_{t-1}))+\mathcal{N}(0,R^{2}\sigma^{2}\mathbf{I}_{d}),$ (7)
where $\mathbf{I}_{d}$ is an identity matrix with size $d\times d$ ($d$ is the model size, i.e., $\bm{\theta}_{t}\in\mathbb{R}^{d}$), the record-level clipping $\mathsf{Clip}_{R}(\cdot)$ restricts the sensitivity when adding/removing one record from the local dataset, and Gaussian noise $\mathcal{N}(0,R^{2}\sigma^{2}\mathbf{I}_{d})$ introduces DP property on $\bm{g}_{t,i}$. Since DP is immune to post-processing, the remaining steps can follow the original LFH without additional privacy costs. This baseline solution DP-LFH achieves record-level DP against an untrusted server but has limitations, leading to poor privacy-utility tradeoff and robustness.
Limitation 1: large aggregated noise. Since each client locally adds DP noise, the overall noise after aggregation is larger than the case of the central setting under the same privacy budget $\epsilon$ since only the server adds noise in the central setting. Therefore, DP-LFH has a poor privacy-utility tradeoff.
@@
where $\bm{g}_{1,i},\bm{g}_{2,i},\cdots,\bm{g}_{t,i}$ are already privatized via local noise. Assume the server samples a subset of clients for aggregation in each iteration. If client $\mathsf{C}_{i}$’s momentum $\bm{m}_{t,i}$ is not selected in the $t$-th iteration, the aggregate is independent of $\bm{g}_{t,i}$. However, in a later iteration (i.e., $\tau>t$), if client $\mathsf{C}_{i}$’s momentum $\bm{m}_{\tau,i}$ is included, it depends on $\bm{g}_{t,i}$ according to (9). Thus, we must account for the privacy cost of $\bm{g}_{t,i}$ in all iterations. Sampling clients offers no privacy amplification, resulting in high privacy costs or low utility.
@@ 4 DP-BREM
To address DP-LFH’s limitations, we propose DP-BREM, a differentially-private LFH variant assuming a trusted server that generates DP noise. DP-BREM maintains robustness of LFH and uses a different privacy accountant (Theorem 1) than DP-SGD. We also provide convergence analysis (Theorem 3) showing minimal deviation from LFH. We further relax the server trust assumption in DP-BREM+ (Section 5) by using secure multiparty computation for secure aggregation and joint noise generation, achieving the same DP and robustness guarantees without a trusted server.
@@ 4.1 Algorithm Design
The illustration of our design is shown in Figure 1, and the algorithm is shown in Algorithm 1, where all clients need to implement local updates (in Line-3), but only a subset of their momentum vectors are aggregated by the server (in Line-4). The details of client updates and server aggregation are described below.
@@
where $\mathsf{Clip}_{R}(\cdot)$ is the clipping function defined in (4), but is used here to bound the sensitivity for DP (refer to DP-SGD discussed in Section 2.1). $\mathcal{D}_{i,t}$ represents a random subset obtained via subsampling from client $\mathsf{C}_{i}$’s dataset. This subsampling is essential to apply privacy amplification, enabling the privacy accountant to derive a tight privacy budget $\epsilon$. Note that the batch size $|\mathcal{D}_{i,t}|$ is random and $\mathbb{E}[|\mathcal{D}_{i,t}|]=p_{i}|\mathcal{D}_{i}|$. Then, the local momentum can be computed by
@@
0:  Initialization $\bm{\theta}_{0}\in\mathbb{R}^{d}$, clipping bounds $R$ and $C$, learning rate $\eta_{t}$ of the global model, total number of iterations $T$, client-level sampling rate $q$, record-level sampling rate $p_{i}$.
@@ 4.2 Privacy Analysis
(Sketch) According to (10), the sensitivity of $\bar{\bm{g}}_{t,i}$ is $\frac{R}{p_{i}|\mathcal{D}_{i}|}$ because each clipped term $\mathsf{Clip}_{R}(\cdot)$ has bounded L2-norm, i.e., $\|\mathsf{Clip}_{R}(\cdot)\|\leqslant R$. Then, due to the recursive representation of local momentum in (11), the sensitivity of $\bm{m}_{t,i}$ is $\frac{R}{p_{i}|\mathcal{D}_{i}|}$. Finally, the client-level clipping $\mathsf{Clip}_{C}(\cdot)$ introduces another upper bound for the sensitivity. Refer to Appendix H for the detailed proof. ∎
@@
The final privacy analysis of DP-BREM is shown in Theorem 1. It presents how to compute the privacy budget $\epsilon$ and privacy parameter $\delta$ when the parameters (such as $T$, $\sigma$, $q$, etc.) of the algorithm are given. We use an advanced privacy accountant tool called Gaussian DP (GDP) [12], then convert it to $(\epsilon,\delta)$-DP. Note that in our privacy analysis, clients can use different record-level sampling rates $p_{i}$, thus different sensitivity $S_{i}$ shown in (14). Therefore, the final privacy budget (denoted by $\epsilon_{i}$) of DP-BREM may be different for different clients, which provides personalized privacy if these parameters are different for each client.
@@
This result is obtained by the composition of multiple iterations and the privacy amplification from sampling. See Appendix A for the detailed proof. ∎
@@ 4.3 Convergence Analysis
After some parameter tuning (the detailed tuning is shown under (B) in Appendix B) of the clipping bounds:
@@
where $|\mathcal{B}|$ is the number of Byzantine clients, $d$ is the dimension of model parameter $\bm{\theta}_{t}$, $\sigma$ is the noise multiplier (for DP) shown in (12), $q$ is the client-level sampling rate shown in Line-2 of Algorithm 1, and $\rho$ is defined in (17). The formal version of (19) is shown in (B) of Appendix B.
@@
(Sketch) Directly bounding $\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{m}_{t}^{*}\|^{2}$ is not easy, thus we utilize the upper bounds of $\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{\mu}\|^{2}$ and $\mathbb{E}\|\bm{\mu}-\bm{m}_{t}^{*}\|^{2}$ to get the final result, where $\bm{\mu}\coloneqq\mathbb{E}[\bm{m}_{t,i}]$ is the expected local momentum (we assume clients’ local momentum are i.i.d.). When upper bounding $\mathbb{E}\|\tilde{\bm{m}}_{t}-\bm{\mu}\|^{2}$, we decompose errors into three types: honest clients’ error (from clipping randomness and bias), Byzantine clients’ error (from perturbation), and DP noise error. Optimizing parameters $C$ and $R$ minimizes the total error. See Appendix B for the detailed proof. ∎
@@
Table 2: Comparison of Convergence Rate
Where to add noise Convergence Rate
LFH [23] None $O(\rho\sqrt{1+|\mathcal{B}|})$
@@ 5.1 Challenges
Secure Aggregation with Verified Inputs (SAVI). The key crypto technique we leverage to achieve the above objectives is SAVI [36], which is a type of protocols that securely aggregate only well-formed inputs. The security goals include both privacy and integrity. Specifically, privacy means that no party should be able to learn anything about the raw input of an honest client, other than what can be learned from the final aggregation result. Integrity means that the output of the protocol returns the correct aggregate of well-formed input, where 1) an input $u$ passes the input integrity check with a public validation predicate $\mathsf{Valid}(\cdot)$ if and only if $\mathsf{Valid}(u)=1$, and 2) the aggregation is correctly computed. An instantiation of the SAVI protocol is EIFFeL [36] (described in Appendix J).
@@
A straightforward approach is to use a semi-honest server, as proposed in [37], to generate DP noise and manage the privacy engine. However, relying on another non-colluding server may be impractical, so we assume only a single server. An alternative is Distributed DP [40], where clients locally generate Gaussian noise. The aggregated noise follows a Gaussian distribution with an enlarged standard deviation, ensuring DP through cryptographic techniques. This method, however, has two limitations: it requires more noise to achieve the same privacy level due to potential collusion among malicious clients, and the robustness is compromised as malicious clients can generate arbitrary local noise.
@@
Overview of DP-BREM+. To achieve secure aggregation with verified inputs and secure Gaussian noise generation under the threat model of a malicious server and malicious minority of clients, our DP-BREM+ 1) leverages an existing SAVI protocol called EIFFeL [36] to achieve secure input validation; and 2) introduces a new protocol to achieve secure noise generation that is compatible with EIFFeL. The idea of jointly generating Gaussian noise in DP-BREM+ is inspired by [35], but our design is based on Shamir’s secret sharing [38] with robust reconstruction by following the design in EIFFeL, thus guarantees security under malicious minority. We present the preliminaries of Shamir’s secret sharing and EIFFeL protocol in Appendix J.